OpenStudy (anonymous):

Algebra II : Find the standard form equation of the parabola with vertex (0,0) and directrix x= 9/2. Can someone please explain how to solve this one? (and solve. We got the answer, x=-18y^2, I just want to make sure its right.

OpenStudy (whpalmer4):

Mmm...the directrix isn't right...

OpenStudy (whpalmer4):

The vertex form equation for a parabola with a horizontal axis of symmetry and vertex at \((h,k)\) is \[(y-k)^2=4p(x-h)\]\(p\) is the distance from vertex to focus and vertex to directrix. Our vertex is at \((0,0)\) so that becomes\[y^2=4px\]\[p=0-\frac{9}{2} = -\frac{9}{2}\]so we have\[y^2=4(-\frac{9}{2})x\]\[ y^2=-18x\] A plot of the parabola is attached.